For the DC circuit shown below, what is the equiva
Answer #1

R2, R4 and R3 are in parallel connection.

equivalent of these three resistance is R' then

1/R' =1/R2 + 1/R4 + 1/R3 = 1/50 + 1/75 + 1/50

R' = 18.75 Ohm

now R' and R1 are in sereis.

Req = R' + R1 = 18.75 + 100 = 118.75 Ohm ........Ans


Current throught battery, I = V / Req =6 /118.75 = 0.0505 A Or 50.5 mA

through R1, I1 = I =50.5 mA

PD across R' = e - I1r1 =6 - (0.050 x 100) = 0.9474 Volt

I2 = 0.9474 / 50 = 0.0189 A   Or 18.9 mA

I3 = 0.9474/50   = 0.0189 A Or 18.9 mA

I4 = 0.9474/75 = 0.0126 A Or 12.6 mA